When using a pencil eraser, you exert a vertical force of 6.00 N at a distance of 2.00 cm from the hardwood-eraser joint. The pencil is 6.00 mm in diameter and is held at an angle of 20.0° to the horizontal. a. By how much does the wood flex perpendicular to its length? b. How much is it compressed lengthwise?
1
Expert's answer
2020-03-10T10:56:37-0400

Given Data

  • The diameter of the pencil is: d=6mm=6×10−3m
  • The vertical force exerted on pencil is: P=6N
  • The angle made by the pencil is: θ=20∘
  • The original length of the pencil is: l=2cm=2×10−2m

The expression to calculate the area is given by,

A=πr2=π(d/2)2

Substitute the values in above equation.

A=π(6×10−32)=2.82×10−5m2

(a)

The expression to calculate the change in length, force acted perpendicular to area using the formula is given by,

E=σ/ξ=((Pp/A)/(Δl/l)); Δl=Pp/A×l/E=(Pcosθ)/A×l/E

Here, σ is stress, ξis strain and the Eis Young's modulus.

Substitute the values in above equation.

Δl=(6cos20∘)/(2.82×10−5)×(2×10−2)/(200×109)=1.999×10−8m

Thus, the wood flex perpendicular to its length is 1.999×10−8m


(b)

The expression to calculate the elongation in pencil using the formula of shear modulus is given by,

G=τ/γ=((P/sinθ)/A)/(Δx/l); Δx=(Psinθ)/A×l/G

Here, τ is shear stress, γ is shear strain.

Substitute the values in above equation.

Δx=(6sin20)/(2.82×10−5)×(2×10−2)/(15×109)=9.7×10−8m

Thus, the compressed length is 9.7×10−8m .